Maple Syrup prices soared for a second year. A combination of warm weather and maple-tree die-off from acid rain again reduced the syrup supply. One farm’s output fell to only 1,200 gallons this year, compared with a normal 4,200. Another’s fell only 30% to 5,000 gallons.

A gallon of syrup sells for $35 at one outlet in Massachusetts, up from $25 a year ago. Some producers charge more. One concern is that the higher prices will scare off food processors and consumers. But one producer noted that "It's a unique product and demand remains firm."

TRUE OR FALSE?

3i. As a result of the decrease in supply of maple syrup, demand is expected to decrease also.

3ii. The statement "It's a unique product and demand remains firm" suggests that demand is price elastic.

3iii. Assuming price increased from $25 to $35, if quantity purchased fell by 11 percent, then the price elasticity of demand is .33 (using midpoint formula).
Answer #1

a) As the supply of the syrup has decease it will increase the price of the good and a higher price will decrease the quantity demand in the market. Quantity demand is different form the demand, the answer is "False". Demand will not decrease but quantity demanded.

b) "False"

the statement that the demand is unique and remains firm means that the demand is inelastic i.e. it will not change due to a change in the price.

c) % change in price = 35 -25 / ((35+25)/2 )

= 10 / (17.5 + 12.5)

= 10 / 30

= 0.33 or 33%.

Demand elasticity = % change in the demand / % change in the price.

= 11 / 33

= 0.33. The statement is "true."